write a system of linear equation whose solution gives the number of each type of room rented. when ten rooms are rented at the rates of $80 and $120 pernight. The total collected of the 10 rooms is $920.
Answer by ptaylor(2198)

Let x=number of rooms rented at $80/night; total collected from these rooms=$80x
And let y=number of rooms rented at $120/night; total collected=$120y
Now we are told that:
$80x+$120y=$920--------------------eq1
And we are also told that:
x+y=10-------------------eq2
We can solve these:
First divide each term in eq1 by 40 and we get:
2x+3y=23 ($ understood)----eq1a
Multiply each term in eq2 by 2 and subtract it from eq1a and we get:
y=3 ----number of $120/night rooms
substitute y=3 into eq2 and we have x=7----number at $80/nite
CK
3*$120+7*80=$920
$360+$560=$920
$920=$920
Hope this helps-----ptaylor